# A Survey on Automatic Credibility Assessment of Textual Credibility Signals in the Era of Large Language Models

### v1

@@ 8.5. Originality and content reuse
In the context of Natural Language Processing, plagiarism detection has become more difficult due to the amount of texts available both in traditional print publications and now online. The challenge lies in detecting not only direct copying but also other forms of plagiarism, such as paraphrasing, translations or using hired writers to produce content. The exploration of originality and credibility signals in text reuse involves a variety of approaches combining NLP techniques with advanced machine learning methods. One prominent strategy is leveraging sentence segmentation, tokenization, and syntactic parsing to dissect the structure of text (Chong et al., 2010). The proposed framework integrates these NLP techniques with trigram similarity measures and dependency relations matching, showing a detection accuracy of 70.53% for plagiarized short excerpts of text. This proves how traditional linguistic tools can be enhanced by statistical models to detect subtle patterns in content reuse. Similarly, trigram similarity alongside language model probability and longest common subsequence methods to deal with plagiarism detection can be adopted (Pal et al., 2023). Use of a Naive Bayes classifier resulted in the same accuracy rate of 70.53%, underscoring the importance of pairing conventional text analysis techniques with machine learning algorithms to capture nuanced instances of content duplication.
@@
Beyond general plagiarism detection, more specialized approaches concentrate themselves on verifying the authenticity of authorship. Enriquez et al. (2023) developed an Authorship Verification model aimed at identifying hired plagiarism, where someone other than the credited author produces the work. By analyzing stylistic features, such as writing patterns and comparing them with the claimed author’s profile, their approach achieved an accuracy of 85%. This sheds light on the possibilities of stylistic analysis as a signal of originality, distinguishing between genuine and outsourced authorship.
The datasets used in these studies on plagiarism detection can be found from a variety of sources and text types, reflecting the different challenges posed by plagiarism. For example, Bin-Habtoor and Zaher (2012) focus on general plagiarism detection systems, analyzing publicly available datasets from academic publications. Chong et al. (2010) utilize a dataset of academic articles to explore plagiarism through techniques like trigram similarity and dependency relations, applying them to short passages to measure detection accuracy. Enriquez et al. (2023) use a dataset of student writing to train their Authorship Verification model, particularly aimed at detecting plagiarism in commissioned work. Elngar et al. (2021) apply their algorithm to a dataset comprising academic and non-academic texts, with an emphasis on lexical, semantic, and syntactic analysis. Lastly, Pal et al. (2023) use a dataset of online content, including articles and essays, to experiment with automatic plagiarism detection using advanced NLP techniques, combining feature extraction with machine learning models. These datasets represent a range of content, allowing for the robust evaluation of plagiarism detection methods across different textual formats.
@@ 8.6. Offensive language
The use of offensive language (and its related phenomena, such as hate speech, abusive or toxic language; see, e.g., (Papcunová et al., 2023) for definitions), particularly in online discourse, often correlates with the low credibility of the content, containing false information, prejudices, biases, and what in general is considered as toxic language. For example, Botella-Gil et al. (Botella-Gil et al., 2024) show that there is a direct association between the credibility of media sources and the presence of hate in online comments, while Bourgeade et al. (Bourgeade et al., 2023) show how fake stories are used to spread hate against immigrants by analyzing racial hoaxes on Twitter. Also some of the most widely used datasets for check-worthy claim detection already include information about the presence of hateful (Hadj Ameur and Aliane, 2021) or harmful (Nakov et al., 2022a) content, showing the importance of adopting a multi-faceted view on the problem of misleading information. Indeed, the use of abusive or toxic language, has also been included in the W3C list of credibility signals (Multiple authors and Hawke, 2019).
@@
Offensive language detection has been extensively investigated within the NLP community for at least a decade, starting from English (Waseem and Hovy, 2016) and then comprising more and more languages (for example the Hate Speech Datasets repository272727https://hatespeechdata.com/ that lists more than one hundred datasets in 25 languages). The creation of resources, which led in turn to the development of several approaches to offensive language detection, has been fostered by the shared tasks on hate or offensive language detection organised throughout the years at SemEval and other evaluation campaigns (Fersini et al., 2018; Bosco et al., 2018; Wiegand et al., 2018; Zampieri et al., 2019, 2020; Pavlopoulos et al., 2021). Among the different approaches implemented for offensive language detection, transformer-based architectures based on RoBERTa (Liu et al., 2019b) and its multilingual variants have proven to be very effective for the task when fine-tuned on offensive data, with F1 scores 0.90 (Zampieri et al., 2020). Nevertheless, the problem of online content moderation is far from being solved given the amount of toxic content still circulating on social media, which suggests that research on the topic still needs to address several understudied aspects of the phenomenon. Current works have identified some research directions worth studying, such as the problem of human label variation when creating datasets for offensive language detection (Leonardelli et al., 2021), the role of annotators’ biases (Sap et al., 2022), the presence of spurious correlations in existing datasets (Wiegand et al., 2019; Ramponi and Tonelli, 2022) and the lack of robustness when classifying data from different online platforms and domains (Salminen et al., 2020). Concerning generative LLMs, recent experiments showed that classifying offensive language using zero- and few-shot learning LLMs yields considerably lower results in comparison to smaller models fine-tuned on the entire training set (Edwards and Camacho-Collados, 2024). Generative LLMs have been alternatively used to create synthetic datasets for hate speech detection, with the goal to address problems such as data decay and privacy concerns related to social media (Casula et al., 2024). However, the fact that major generative LLMs have been adjusted through so-called alignment to avoid generating hateful content may represent an obstacle for future research in this direction.
@@ 8.7. Non-NLP research on credibility signals
While the goal of this survey is to give an overview of the credibility assessment and credibility signals from the perspective of Natural Language Processing, additional credibility signals that can be extracted from ancillary elements are not of lesser relevance. These non-NLP credibility signals can range from the multimedia attached or linked to the textual information (Zade et al., 2023; Hansen et al., 2020; Hamborg et al., 2017), to other cues based on different metrics, such as the impact of an author (Briscoe et al., 2013; Jahanbakhsh et al., 2022; Zhao et al., 2016), or even the acceptance of the engaged public (İŞ and TUNCER, 2018; Kondo and Abe, 2023; Lin et al., 2016). For example, it can be expected that a fake video would unlikely be a part of credible content; likewise, a biased author with an untrustworthy historic record could lead to the same outcome, the same as the refutation of an accepted authority in the field.
Several taxonomies can be used to categorize non-NLP credibility signals, for example, a distinction can be made on the signals coming from different content such as complementary videos or images, or external ones, like author or source information (Ye and Skiena, 2019; Patricia Aires et al., 2019; Aires et al., 2020a). A general classification of non-NLP credibility signals is summarized in Table 10 with corresponding signals’ examples and references.
@@
As presented, a variety of factors could be leveraged in order to perform a credibility assessment, and the potential usage of them is a decision to make depending on the efficiency and the availability of the signals in the experiment. It is noticeable that the credibility signals are from a large variety of sources and formats, and some of them have been presenting good research results recently (Ciobanu et al., 2014; Ye and Skiena, 2019; Riegelsberger et al., 2005). Additionally, it is worthy of noticing that this pattern is also shown through the extensive variety of credibility signals present in the text, where a large variety of credibility signals and much active presence in the research field can also be found, due to the big potential of the NLP techniques in this regard.
Lastly, a mention of the possibility of applying both textual and non-textual signals combined in the process of credibility assessment should be given, as multiple studies support these types of methodologies with good experimental results (Briscoe et al., 2013; Jahanbakhsh et al., 2022; Yan et al., 2024). For example, one common approach to tackle low credible content detection is the usage of fake news propagation graphs (Solanki et al., 2021; Chua et al., 2023; Zade et al., 2023), where the social activity of the sharing and propagation of a piece of news in a social network is the subject of study. There exists another aspect of interest in these studies where there is also a common practice of combining both the user and community engagement together with the actual content (including using NLP techniques to process the text) to detect the trustworthiness of the piece of news in particular (İŞ and TUNCER, 2018; Lin et al., 2016; Zhao et al., 2016).
The practice of applying both NLP and non-NLP based procedures, jointly with the notable variety of credibility signals (Rijo and Waldzus, 2023; Ji et al., 2023b; O’Grady, 2006) shows the interest and the open possibilities to aid in the field of exploiting these findings. The usage of one or the other, or both, will be dependent on the specific case of usage.
@@ 9. Challenges and open problems
Following the analysed research works as well as our own experience in the area of credibility assessment and credibility signals, we follow up with an overarching and orthogonal (to credibility assessment and categories of credibility signals described so far) discussion of challenges and open problems.
@@ 9.1. Adoption and potential of generative LLMs
Such models offer several opportunities to address challenges related to the automatic detection of credibility signals. One of the key advantages of using prompting with LLMs is the flexibility in adapting a single foundational model to handle multiple subtasks associated with credibility assessment. With a carefully designed framework, LLMs can be guided to focus on different aspects of content analysis, such as detecting persuasion techniques, evaluating the veracity of claims, identifying potential bias, and recognizing patterns of misinformation. This flexibility reduces the need to develop and fine-tune separate models for each task, allowing practitioners to use the same model across various credibility-related tasks. In fact, a promising research direction is to explore multi-task learning, as in verifying if the capacity of performing certain credibility-assessment tasks can aid in other related ones (e.g., persuasion and bias).
@@
Moreover, the capability of learning with zero/few examples is particularly valuable for tasks where domain-specific data is scarce or constantly changing, as is the case with credibility assessment. An enormous amount of human effort is required to curate high-quality annotated datasets for the different subtasks involved in assessing credibility. Specially since tasks such as labeling misinformation or identifying biased content often demand the expertise of domain specialists such as fact-checkers, journalists and social scientists. Adding to this challenge, credibility indicators can be highly context-dependent, varying across cultural and temporal dimensions. In this context, LLMs offer a more scalable approach by drawing on vast amounts of unsupervised pretraining data, and by adjusting to specific end-tasks through careful prompting strategies, which require far less human effort than manual data labeling. As an example, in Leite et al. (2024), a generative LLM was employed to predict different credibility signals present in textual content without using any training data (i.e., in a zero-shot setting).
@@
Finally, the generative capabilities of large language models can be leveraged to produce more interpretable282828Here, the concept of interpretability differs from explainability, which is often used in the field of machine learning to refer to specific methods to analyse how intermediate states of the model lead to certain outcomes (Burkart and Huber, 2021). outputs, which is crucial for subject-matter experts that may leverage the model’s predictions. Instead of providing only binary or scalar outputs (e.g., true/false, misinformation/non-misinformation) as in usual classification tasks, generative LLMs can produce detailed explanations or summaries that can highlight the reasoning behind their predictions. This transparency allows human experts to critically assess the model’s outputs, cross-check them with external information, and ensure that (i) incorrect predictions (in this context, known as model hallucinations (Ji et al., 2023a)) are properly mitigated, and (ii) any credibility assessment aligns with the context of the content being investigated. This property of interpretable outputs can significantly increase trust in model-driven decisions and reduce the likelihood of over-reliance on machine predictions, ensuring that humans remain in the loop for final judgments.

### v2

@@ 8.4. Originality and content reuse
In the context of NLP, plagiarism detection has become more difficult due to the amount of texts available both in traditional print publications and now online. The challenge lies in detecting not only direct copying but also other forms of plagiarism, such as paraphrasing, translations or using hired writers to produce content. The exploration of originality and credibility signals in text reuse involves a variety of approaches combining NLP techniques with advanced machine learning methods. One prominent strategy is leveraging sentence segmentation, tokenization, and syntactic parsing to dissect the structure of text (Chong et al., 2010). The proposed framework integrates these NLP techniques with trigram similarity measures and dependency relations matching, showing a detection accuracy of 70.53% for plagiarized short excerpts of text. This proves how traditional linguistic tools can be enhanced by statistical models to detect subtle patterns in content reuse. Similarly, trigram similarity alongside language model probability and longest common subsequence methods to deal with plagiarism detection can be adopted (Pal et al., 2023). Use of a Naïve Bayes classifier resulted in the same accuracy rate of 70.53%, underscoring the importance of pairing conventional text analysis techniques with machine learning algorithms to capture nuanced instances of content duplication.
@@
Beyond general plagiarism detection, more specialized approaches concentrate themselves on verifying the authenticity of authorship. Enriquez et al. (2023) developed an Authorship Verification model aimed at identifying hired plagiarism, where someone other than the credited author produces the work. By analysing stylistic features, such as writing patterns and comparing them with the claimed author’s profile, their approach achieved an accuracy of 85%. This sheds light on the possibilities of stylistic analysis as a signal of originality, distinguishing between genuine and outsourced authorship.
The datasets used in these studies on plagiarism detection can be found from a variety of sources and text types, reflecting the different challenges posed by plagiarism. For example, Bin-Habtoor and Zaher (2012) focus on general plagiarism detection systems, analysing publicly available datasets from academic publications. Chong et al. (2010) utilize a dataset of academic articles to explore plagiarism through techniques like trigram similarity and dependency relations, applying them to short passages to measure detection accuracy. Enriquez et al. (2023) use a dataset of student writing to train their Authorship Verification model, particularly aimed at detecting plagiarism in commissioned work. Elngar et al. (2021) apply their algorithm to a dataset comprising academic and non-academic texts, with an emphasis on lexical, semantic, and syntactic analysis. Lastly, Pal et al. (2023) use a dataset of online content, including articles and essays, to experiment with automatic plagiarism detection using advanced NLP techniques, combining feature extraction with machine learning models. These datasets represent a range of content, allowing for the robust evaluation of plagiarism detection methods across different textual formats.
@@ 8.5. Offensive language
The use of offensive language (and its related phenomena, such as hate speech, abusive or toxic language; see, e.g., (Papcunová et al., 2023) for definitions), particularly in online discourse, often correlates with the low credibility of the content, containing false information, prejudices, biases, and what in general is considered as toxic language. For example, Botella-Gil et al. (2024) show that there is a direct association between the credibility of media sources and the presence of hate in online comments, while Bourgeade et al. (2023) show how fake stories are used to spread hate against immigrants by analysing racial hoaxes on Twitter. Also some of the most widely used datasets for check-worthy claim detection already include information about the presence of hateful (Hadj Ameur and Aliane, 2021) or harmful (Nakov et al., 2022a) content, showing the importance of adopting a multi-faceted view on the problem of misleading information. Indeed, the use of abusive or toxic language, has also been included in the W3C list of credibility signals (Multiple authors and Hawke, 2019).
@@
Offensive language detection has been extensively investigated within the NLP community for at least a decade, starting from English (Waseem and Hovy, 2016) and then comprising more and more languages (for example the Hate Speech Datasets repository262626https://hatespeechdata.com/ that lists more than one hundred datasets in 25 languages). The creation of resources, which led in turn to the development of several approaches to offensive language detection, has been fostered by the shared tasks on hate or offensive language detection organised throughout the years at SemEval and other data challenges (Fersini et al., 2018; Bosco et al., 2018; Wiegand et al., 2018; Zampieri et al., 2019, 2020; Pavlopoulos et al., 2021). Among the different approaches implemented for offensive language detection, transformer-based architectures based on RoBERTa (Liu et al., 2019b) and its multilingual variants have proven to be very effective for the task when fine-tuned on offensive data, with F1 scores 0.90 (Zampieri et al., 2020). Nevertheless, the problem of online content moderation is far from being solved given the amount of toxic content still circulating on social media, which suggests that research on the topic still needs to address several understudied aspects of the phenomenon. Current works have identified some research directions worth studying, such as the problem of human label variation when creating datasets for offensive language detection (Leonardelli et al., 2021), the role of annotators’ biases (Sap et al., 2022), the presence of spurious correlations in existing datasets (Wiegand et al., 2019; Ramponi and Tonelli, 2022) and the lack of robustness when classifying data from different online platforms and domains (Salminen et al., 2020). Concerning generative LLMs, recent experiments showed that classifying offensive language using zero- and few-shot learning LLMs yields considerably lower results in comparison to smaller models fine-tuned on the entire training set (Edwards and Camacho-Collados, 2024). Generative LLMs have been alternatively used to create synthetic datasets for hate speech detection, with the goal to address problems such as data decay and privacy concerns related to social media (Casula et al., 2024). Similarly, LLMs have also been employed to generate functional tests for a fine-grained evaluation of hate speech detection (Jin et al., 2024). However, the fact that major generative LLMs have been adjusted through so-called alignment to avoid generating hateful content may represent an obstacle for future research in this direction.
8.6. Machine-generated text
The task of machine-generated text detection (also called synthetic or neural text detection or authorship attribution (Uchendu et al., 2023)) has already been researched since about 2018 when the arrival of GPT-1 and transformer architecture have enabled generation of reasonably coherent texts(Uchendu et al., 2023). However, it has been only with the arrival of generative LLMs that the task gained upon practical importance due to increased fidelity and quality of the machine-generated texts in English (Uchendu et al., 2021), but also in many other languages (Macko et al., 2023; Wang et al., 2024b; Macko et al., 2025) and due to low costs of such generation. Since then, there has been recognised potential of the machine-generated texts to be misused for influence operations (Goldstein et al., 2023), disinformation (Buchanan et al., 2021; Vykopal et al., 2024; Lucas et al., 2023; Zugecova et al., 2025), spam or unethical authorship (Crothers et al., 2023). Although there have been voices tempering down fears of massive misuse of LLMs for disinformation generation (Simon et al., 2023), actual field studies are missing and the rapid advancement of the new models makes the misuse of LLMs to generate and/or amplify disinformation increasingly easier. Nevertheless, when considering using machine-generated text detection as a credibility signal, it has to be noted that there are also many benign and legitimate uses (e.g., machine translation, use of LLMs to improve stylistics or grammar, etc.), so the context of such a use needs to be taken into account as well.
Due to increased interest in the topic, there have been lately several new datasets released extending the task to a range of languages and domains (Macko et al., 2023; Wang et al., 2024b; Macko et al., 2025), often supported by specialized data challenges and tasks, such as SemEval-2024 Task 8 (Wang et al., 2024a). As to the detection methods and models, these range from statistical, e.g., Binoculars (Hans et al., 2024), Fast-DetectGPT (Bao et al., 2024), etc. to fine-tuned language models, such as Longformer (Beltagy et al., 2020; Li et al., 2024a), RoBERTa (Solaiman et al., 2019) or MDeBERTa (Macko et al., 2023), but also increasingly including fine-tuned LLMs or their combinations (Spiegel and Macko, 2024).
@@ 8.7. Non-NLP research on credibility signals
While the goal of this survey is to give an overview of the credibility assessment and credibility signals from the perspective of Natural Language Processing (NLP), additional credibility signals that can be extracted from ancillary elements are not of lesser relevance. These non-NLP credibility signals can range from the multimedia attached or linked to the textual information (Zade et al., 2023; Hansen et al., 2020; Hamborg et al., 2017), to other contextual cues, such as the impact of an author (Briscoe et al., 2013; Jahanbakhsh et al., 2022; Zhao et al., 2016), or even the acceptance and feedback provided by the engaged public (İŞ and TUNCER, 2018; Kondo and Abe, 2023; Lin et al., 2016). For example, it can be expected that a fake video would unlikely be a part of credible content; likewise, a biased author with an untrustworthy historic record could lead to the same outcome.
Several taxonomies can be used to categorize non-NLP credibility signals, for example, a distinction can be made on the signals coming from different content modalities, such as complementary videos or images, or contextual entities, like author or source information (Ye and Skiena, 2019; Patricia Aires et al., 2019; Aires et al., 2020a). A general classification of non-NLP credibility signals is summarised in Table 10 with corresponding signals’ examples and references. It is noticeable that the credibility signals are from a large variety of sources and formats, and some of them have been presenting good research results recently (Ciobanu et al., 2014; Ye and Skiena, 2019; Riegelsberger et al., 2005).
@@
Studies have shown that combining textual analysis with multimedia analysis leads to significant improvements in credibility assessment as well as misinformation detection (Choudhary et al., 2024; Alam et al., 2022a; Hangloo and Arora, 2022; Briscoe et al., 2013; Jahanbakhsh et al., 2022). This combination is typically achieved through ensemble models or architectures that unify the outputs of various sub-models focused on different media types. A common method involves using transformer-based models such as BERT or GPT for text analysis alongside convolutional neural networks (CNNs) for image or video analysis. These models work in tandem, the text models identifying linguistic signals of credibility while the CNNs assess visual features like image authenticity, quality, and relevance. Additional layers of analysis might focus on audio quality, transcription accuracy, or even the relationship between the text and accompanying media. For example, a multimodal model might analyse a tweet’s text for sentiment and coherence, while also assessing the attached image’s metadata and authenticity. If the image shows signs of manipulation or does not align with the textual claim, the credibility for that piece of content would be lowered. Another approach is the usage of propagation graphs (Solanki et al., 2021; Chua et al., 2023; Zade et al., 2023), where the social activity of the sharing and propagation of a piece of news in a social network is the subject of study. Finally, some works combine both the user and community engagement together with the actual content (including using NLP techniques to process the text) to detect the trustworthiness of the piece of news in particular (İŞ and TUNCER, 2018; Lin et al., 2016; Zhao et al., 2016).
At the same time, studies in multimodal credibility assessment consistently show that while non-textual cues (e.g., image authenticity, video manipulation detection) significantly enhance the detection process, they are most effective when paired with deep text analysis. For instance, a recent work combines user engagement patterns with textual signals to identify low credibility content (Qureshi and Malick, 2023). Another promising approach is the integration of social media metadata (such as virality metrics and user behaviour) alongside textual content to detect the spread of such low-credible content (Zade et al., 2023; Spezzano et al., 2021).
@@ 9. Challenges and open problems
Following the analysed research works as well as our own experience in the area of credibility assessment and credibility signal detection, we follow up with an overarching and orthogonal (to credibility assessment and categories of credibility signals described so far) discussion of challenges and open problems.
9.1. Research fragmentation and need for integration
The overarching conclusion of this literature survey is the high degree of fragmentation across the field of credibility assessment research. This issue became apparent throughout the entire survey process – beginning with the analysis of existing taxonomies/lists of credibility signals and our unified categorization (Section 2), continuing with the review of research on automatic credibility assessment (Section 4), and extending to approaches for detecting individual credibility signals (Sections 5–8).
This fragmentation has several important consequences. First, there is a lack of standardized taxonomies for credibility signals. The most prominent attempt so far – an extensive list of approximately 200 crowdsourced credibility signals proposed by the W3C Credible Web Community Group (Multiple authors and Hawke, 2019) – remains incomplete and not fully categorized, limiting its practical utility for guiding research and tool development.
Second, there is a significant disconnection within the NLP research community itself. Research on automatic credibility assessment is often conducted in isolation from work on the detection of individual credibility signals. As a result, even highly advanced methods for detecting certain signals (e.g., persuasion techniques) are rarely integrated into broader credibility assessment solutions. Consequently, little is known about how such signals may contribute to predicting overall content credibility in practice.
Third, researchers working on individual credibility signals frequently appear unaware of related efforts in neighbouring categories, despite these signals often being identified under the shared conceptual umbrella of credibility. Stronger interconnection between these lines of research could unlock mutual benefits – for instance, the creation of shared, curated datasets annotated for both individual signals and overall credibility (also building on the first efforts in this direction by (Zhang et al., 2018; Bhuiyan et al., 2020), as discussed in Section 4.4). This would not only increase annotation quality and sample size but also enable the study of correlations between different signals and support the development of multitask learning models – techniques currently underutilized due to a lack of such integrated resources.
Interestingly, although research efforts remain isolated in their own silos built around the single NLP task, some end-user-oriented tools already incorporate multiple credibility signals in practice. Notable examples include the Tanbih system272727https://tanbih.org/ (Zhang et al., 2019) and the Verification Plugin282828https://www.veraai.eu/category/verification-plugin.
To illustrate the potential benefits of detecting and aggregating multiple credibility signals (representing individual categories systematically covered by this survey) into a unified credibility label/score, we provide two illustrative examples: one depicting a low-credibility social media post (Figure 7) and another representing a high-credibility example (Figure 8). These examples clearly showcase an untapped potential of synergy between detection of diverse advanced credibility signals, their explanations, as well as aggregation. To achieve such desired state, we advocate for significantly stronger integration of current research efforts.
@@ 9.2. Adoption and potential of generative LLMs
As we have discussed in the sections on individual credibility signals (see Sections 5, 6, 7, but also 8), the uptake of LLMs differs across the signals. In some cases, first such uses appeared only in late 2024 and beginning of 2025, such as in the case of previously fact-checked claims (Section 7). Nevertheless, their prevalence gradually increases. However, approaches using LLMs for credibility assessment (see Section 4) are still rare with one such notable exception being the work of Leite et al. (2024). Consequently, there are still several unexplored or underexplored opportunities how such models could address challenges related to the automatic detection of credibility signals.
@@
One of the key advantages of using prompting with LLMs is the flexibility in adapting a single foundational model to handle multiple subtasks associated with credibility assessment. With a carefully designed framework, LLMs can be guided to focus on different aspects of content analysis, such as detecting persuasion techniques, evaluating the veracity of claims, identifying potential bias, and recognizing patterns of misinformation. This flexibility reduces the need to develop and fine-tune separate models for each task, allowing practitioners to use the same model across various credibility-related tasks. In fact, a promising research direction is to explore multi-task learning, as in verifying if the capacity of performing certain credibility-assessment tasks can aid in other related ones (e.g., persuasion and bias).
@@
Moreover, the capability of learning with zero/few examples is particularly valuable for tasks where domain-specific data is scarce or constantly changing, as is the case with credibility assessment. An enormous amount of human effort is required to curate high-quality annotated datasets for the different subtasks involved in assessing credibility. Specially since labelling most of credibility signals (such as biased content) often demand the expertise of domain specialists such as fact-checkers, journalists and social scientists. Adding to this challenge, credibility indicators can be highly context-dependent, varying across cultural and temporal dimensions. In this context, LLMs offer a more scalable approach by drawing on vast amounts of unsupervised pretraining data, and by adjusting to specific end-tasks through careful prompting strategies, which require far less human effort than manual data labelling. As an example, in Leite et al. (2024), a generative LLM was employed to predict different credibility signals present in textual content without using any training data (i.e., in a zero-shot setting).
@@
Finally, the generative capabilities of large language models can be leveraged to produce more explainable and interpretable292929Here, the concept of interpretability differs from explainability, which is often used in the field of machine learning to refer to specific methods to analyse how intermediate states of the model lead to certain outcomes (Burkart and Huber, 2021). outputs, which is crucial for subject-matter experts that may leverage the model’s predictions. Instead of providing only binary or scalar outputs (e.g., true/false, misinformation/non-misinformation, credible/non-credible) as in usual classification tasks, generative LLMs can produce detailed explanations or summaries that can highlight the reasoning behind their predictions. This transparency allows human experts to critically assess the model’s outputs, cross-check them with external information, and ensure that (i) incorrect predictions (in this context, known as model hallucinations (Ji et al., 2023a)) are properly mitigated, and (ii) any credibility assessment aligns with the context of the content being investigated. This property of interpretable outputs can significantly increase trust in model-driven decisions and reduce the likelihood of over-reliance on machine predictions, ensuring that humans remain in the loop for final judgments.
@@
Finally, as highlighted in Section 5, implicit biases present in various LLMs can make them unreliable in when assessing external biases or other credibility signals (Lunardi et al., 2024; Bang et al., 2024). Efficient debiasing strategies, evaluation metrics for bias detection and alignment strategies must be implemented before such models can be applied to real-world tasks (Lin et al., 2024a; Mohanty, 2025). Factual correctness of LLMs is also vulnerable to slight contextual shifts and hallucinations (Augenstein et al., 2024), and improvements in source attribution, domain-specific adaptation and reasoning capabilities are cornerstone tasks in this direction that need to be addressed.